OpenStudy (anonymous):

If B = 140◦, a = 64, b = 12, solve the triangle. (If there is no solution, give −1 as your answer.) a) Find A. (If there is more than one solution, give the larger one.) b) Find the corresponding angle C c) Find the corresponding side c If B = 140◦, a = 64, b = 12, solve the triangle. (If there is no solution, give −1 as your answer.) a) Find A. (If there is more than one solution, give the larger one.) b) Find the corresponding angle C c) Find the corresponding side c @Mathematics

OpenStudy (mertsj):

sin140/12=sinA/64 sinA = 64(sin140)/12 sinA = 3.43 This is impossible since the sin cannot be greater than 1. Therefore there is no solution so give -1 as your answer.

OpenStudy (anonymous):

for part b or c?